OpenStudy (anonymous):

Complete the square: y = (x+6)(2x-5)

OpenStudy (anonymous):

first foil, \[x(2x)+x(-5)+6(2x)+6(-5)\] \[2x^2-5x+12x-30\]

OpenStudy (anonymous):

\[2x^2+7x-30\]

OpenStudy (anonymous):

so 2x^2+ 7x - 30 which will take to: 2(x^2 +7/2x) - 30?

OpenStudy (anonymous):

I just don't know what to do from here.

OpenStudy (anonymous):

that would work yes

OpenStudy (anonymous):

alright so first take the coefficient of the x value and divide it in half

OpenStudy (anonymous):

7/4?

OpenStudy (anonymous):

\[\frac{7}{2}*\frac{1}{2}\]

OpenStudy (anonymous):

yes now square that

OpenStudy (anonymous):

Oh god. This will be a big number. 49/16?

OpenStudy (anonymous):

yep so now thats your second square.. aka b in (a+b)^2 so now if you add it into the parenthesis you have to subtract it on the outside so that you don't manipulate the equation . In otherwords \[2(x^2+\frac{7}{2}x+\frac{49}{16})-30-\frac{49}{16}=0\]

OpenStudy (anonymous):

so now you have a perfect square within the parenthesis \[2(x+\frac{7}{4})^2-30-\frac{49}{16}=0\]

OpenStudy (anonymous):

to simplify more you can turn the 30 into 16th's and add them together

OpenStudy (anonymous):

don't you have to multiply 49/16 by 2?

OpenStudy (anonymous):

yes you want to multiply that number by 2

OpenStudy (anonymous):

it should look like this \[2[(x+\frac{7}{4})^2-\frac{49}{16}]-30\]

OpenStudy (anonymous):

so 90/16? then from there it's - 480/16 - 90/16 2(x + 7/2)^2 - 285/8

OpenStudy (anonymous):

oh wait.

OpenStudy (anonymous):

but then arent you multiplying the x by 2 which will make it 2x?

OpenStudy (anonymous):

meh i'd leave it out if you're teacher doesn't mind

OpenStudy (anonymous):

i'd just simplify the outside of the parenthesis and leave it as 2( )

OpenStudy (anonymous):

Alright, i'll just wait for a solid answer. Thanks for the help :D